OpenStudy (anonymous):

A store mixes Kentucky bluegrass worth $12 per pound and chewings fescue worth $15 per pound. The mixture is to sell for $13 per pound. Find how much of each should be used to make a 228 pound mixture. How many pounds of the Kentucky bluegrass should be used? How many pounds of chewings fescue should be used?

OpenStudy (ganpat):

u can try this... Let x = Kentucky bluegrass and y = chewings fescue we can get 2 equations.. x + y = 228 12x + 13y = 15 (228) can u solve further ??

OpenStudy (anonymous):

yes thank you

OpenStudy (ganpat):

w/c :)